If 2/3 of the teachers in a school own a house, what is the ratio of non-homeowners to homeowners?
a. 2:5
b. 1:2
c. 3:5
d. 3:2

Answer by jim_thompson5910(35256)

Let's say there are 30 teachers. Two-thirds of them is (2/3)*30 = (2*30)/3 = 60/3 = 20. So 20 teachers own a house. The remaining 30 - 20 = 10 teachers don't own a house.


ratio of non-homeowners to homeowners = (non-owners) : (owners)

ratio of non-homeowners to homeowners = 10:20

ratio of non-homeowners to homeowners = 1:2


Which effectively says "there are twice as many owners as there are non-owners"